OpenStudy (anonymous):

In right triangle ABC, CD is the altitude to hypotenuse AB. if CD = 6 and the ratio of AD is 1:5, determine and state the length of BD

OpenStudy (anonymous):



OpenStudy (anonymous):

Use the formula: (CD)^2 - AD x DB Cal AD = x and BD = y. you have 2 equations to find x. y = 5x x*y = 6^2 = 36 -> x*(5x) = 36 -> 5x^2 = 36 x^2 = 36/5 -> x = 6/V5 = 6*V5/5 y = 5x = 6V5

OpenStudy (anonymous):

Correction: (CD)^2 = AD x DB